OpenStudy (anonymous):

diff eq..determine the form of a particular solution for the given nonhomogeneous equation y''+2y'=3-4t

OpenStudy (amistre64):

yp = some educated guess yp=At + B ; if these are not already a linear solution to the homogenous part ...

OpenStudy (anonymous):

for fundamental solution set, i have y1=1, y2=e^(-2t) f(t)=3-4t polynomial method so yp=-At+B...check to see if any matches the fund.set. it does so... yp=-At^2+Bt find y'p and y''p and plug in. now solving for A and B is where i'm stuck

OpenStudy (amistre64):

then lets work thru it so I can catch up :)

OpenStudy (anonymous):

ok thx

OpenStudy (amistre64):

r^2+2r=0 r=0,-2 right?

OpenStudy (anonymous):

at this point, it's just algebra which apparently, i'm very weak at

OpenStudy (anonymous):

right

OpenStudy (amistre64):

\[y=c_1e^{0x}+c_2e^{-2x}+y_p\] \[y=c_1+c_2e^{-2x}+y_p\]

OpenStudy (amistre64):

yp = At + B ; but we already have a constant so we need to up the ante yp = At^2 + Bt ; and derive

OpenStudy (amistre64):

yp = At^2 + Bt yp'= 2At + B yp''= 2A agreed?

OpenStudy (amistre64):

yp''+2yp'=3-4t (2A) + 2(2At + B)=3-4t im going to do this in lowercase, just easier to type for me 2a +4at +2b = 3 -4t (4a)t = -4t ; a=-1 (2a+2b) = 3 -2+2b = 3 ; b=5/2

OpenStudy (anonymous):

ah ok, i almost did that to solve for a but wasn't sure if it was correct, thank you

OpenStudy (amistre64):

at^2 + bt = -t^2+ 5/2 t\[\] \[y=c_1+c_2e^{-2x}-t^2+\frac 52 t \]